OpenStudy (anonymous):

A ball of moist clay falls 10.0 m to the ground. It is in contact with the ground for 21.0 ms before stopping. (a) What is the average acceleration of the clay during the time it is in contact with the ground? (Treat the ball as a particle.) (b) Is the average acceleration up or down?

OpenStudy (amistre64):

change in velocity with respect to the change in time

OpenStudy (amistre64):

we can determine the velocity with energy equations since the total kinetic energy at the bottom is equal to the total potential energy that it started with mgh = 1/2 mv^2 2gh = v^2